OpenStudy (anonymous):

the volume of a sphere can be give by v-4.18879r^3. you have to design a spherical cotainer that will hold a volume of 55 cubic inches. what should the radius be ?

OpenStudy (anonymous):

If we solve the given equation for r in terms of v, we get \[r = \sqrt[3]{v/4.18879}.\]

OpenStudy (anonymous):

Plug 55 into that formula , get a numerical answer and you're done.

OpenStudy (anonymous):

it didnt come out too one of my answer choices .

OpenStudy (anonymous):

I got 13.13 inches.

OpenStudy (anonymous):

If that is incorrect (I suspect it is), then you have bad data. The correct formula for the volume of a sphere is \[V = (4/3)pir^3.\]

OpenStudy (anonymous):

Lets see what happens if we work with this, going through the same process again.

OpenStudy (anonymous):

\[r = \sqrt[3]{3V/(4*\pi)}\].

OpenStudy (anonymous):

I still get 13.13 inches.

OpenStudy (anonymous):

Forgot the cube root of that, which is 2.36 in.